# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ 3.2 Option Pricing: A Novel Analytical Approximation
The European call option price at time $t$ is given by the conditional expectation $C_{t}=e^{-r\left(T-t\right)}\mathbb{E}_{t}^{\mathbb{Q}}\left[\max\left(S_{T}-K,0\right)\right]$, where $T$ is the maturity date, $K$ is the strike price, and $S_{T}$ is the terminal price of underlying asset. Let $M=T-t$ be the number of periods to maturity, it is well-known that an affine structure of the MGF of future cumulative returns,
$g_{t,M}(s)=\mathbb{E}_{t}^{\mathbb{Q}}\left[\exp\left(s\sum_{i=1}^{M}R_{t+i}\right)\right],\quad s\in\mathbb{R}$ (14)
@@ 3.2.1 MGF with Predetermined Path
Suppose that Assumptions 1 and 2 hold and that $\log\eta_{t}$ take the predetermined path, $\bar{\boldsymbol{\eta}}_{t,M}$. Then the MGF for future cumulative returns under $\mathbb{Q}$ has the affine form,
@@ 3.2.2 Option Pricing with Stochastic Variance Risk Aversion
We now turn to the more challenging problem where $\boldsymbol{\eta}_{t,M}$ is stochastic. The lack of an affine structure thwarts the standard approach to obtaining closed-form option pricing formulae, and it is common to resort to simulation methods in this case.
@@
Suppose that Assumptions 1-3 hold. The approximate model-implied MGF, based on a second-order Taylor expansion, is given by
$\hat{g}_{t,M}(s)=g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M}^{e})\left[1+\frac{1}{2}{\rm tr}\{H_{t,M}(s)\Sigma_{M}\}\right],$ (15)
@@
Theorem 3 clarifies how random variation in $\eta_{t}$ impacts the MGF. The core principle behind the proposed approximation method is akin to perturbation methods used to find approximate solutions to Schrödinger equations for which standard methods are not applicable. Perturbation methods also starts from an exact solution of a simpler, related problem, from which an approximate solution to the actual problem is deduced. Our approximation is based on a Taylor expansion about an $M$-dimensional vector, where the conditional moments of the terms in the expansion are deduced from Assumption 3. Next, we evaluate the accuracy of the approximation.
Figure 3: Risk-neutral densities of (six months) cumulative return, for different initial values of variance risk ratio, $\eta_{0}=0.70$ (left panel), $\eta_{0}=1.11$ (middle panel), and $\eta_{0}=1.70$ (right panel). The blue solid line represents the true density from 1 million simulations, the red dashed line is the (2nd-order) approximate density by DHNG, the gray dash-dotted line is the auxiliary model based on $\bar{\boldsymbol{\eta}}_{t,M}^{e}$ (1st-order), and the yellow dotted line is the (0th-order) density by CHNG with constant $\eta_{t}$.
@@ 3.3 Accuracy of Analytical Approximation
The option pricing formula, (16), involves a second-order approximation that accounts for the first two conditional moments of $\boldsymbol{\eta}_{t,M}$. [^8] If the distribution of $\boldsymbol{\varepsilon}_{t,M}$ is symmetric, (e.g. Gaussian), then it is an approximation to third-order. A first-order approximation is equivalent to assuming that $\eta_{t}$ will take the path of its conditional moment, $\bar{\boldsymbol{\eta}}_{t,M}^{e}$. Thus, $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{e})$ is the MGF for the first-order approximation, which we previously labeled “Auxiliary”, because it was used as an intermediate step towards our preferred approximation, $\hat{g}_{t,M}(\cdot)$. Taking $\eta_{t}=\eta_{0}$ to be constant, $\bar{\boldsymbol{\eta}}_{t,M}^{c}=(\log\eta_{0},\ldots,\log\eta_{0})^{\prime}$, is a characteristic of CHNG, and this case can be interpreted as the zero-order “approximation”.
We first consider the risk-neutral density of cumulative returns. Figure 3 presents risk-neutral densities for cumulative returns over six months (blue solid lines) for three initial values, $\eta_{0}=0.70$ (left panel), $\eta_{0}=1.11$ (middle panel), and $\eta_{0}=1.70$ (right panel). The data generation process is based on the structure in Theorem 1, using the parameter estimates from our empirical analysis of option prices with $\log\eta_{t}\sim\operatorname{AR}(1)$, normally distributed innovations, and $h_{1}$ initialized at its unconditional mean. In this design, the exponential of unconditional mean, $\exp\left(\mathbb{E}\log\eta_{t}\right)=\exp\left(\zeta\right)$, is used as the initial value in the middle panel. The true densities are obtained from 1 million simulated time-series.
The red dashed lines are the risk-neutral densities of DHNG, i.e. those implied by $\hat{g}_{t,M}(\cdot)$, the gray dot-dashed lines are based on the auxiliary, $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{e})$, and the yellow dotted line is the risk-neutral density for CHNG, i.e. $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{c})$. DHNG, which is based on the second-order approximation of Theorem 3, is accurate, whereas the Auxiliary structure fails to match the upper tail of the densities. The CHNG is the worst approximation of the true density in all cases. It is tied with Auxiliary in the middle panel, because the two are identical when $\log\eta_{0}$ is set to have its unconditional mean $\zeta$.
@@
Figure 4: Annualized variance, $\frac{252}{M}{\rm Var}_{t}^{\mathbb{Q}}\left(\sum_{i=1}^{M}R_{t+i}\right)\times 100\%$, skewness, and kurtosis of cumulative returns over $M$ days, are shown with three different initial variance risk ratios, $\eta_{0}=0.70$ (left panels), $\eta_{0}=1.11$ (middle panels), and $\eta_{0}=1.70$ (right panels). Blue lines represent true population moments, red dashed lines are those of DHNG, yellow dotted lines are those for CHNG, and the gray dash-dotted lines are those for the auxiliary “model”. Population moments were obtained from $10^{6}$ simulated trajectories, using a design based on the parameter estimates from obtained with option prices.
@@
The risk-neutral $k$-th moment for $k\in\mathbb{N}$ of $M$-th ahead cumulative returns can be expressed as
@@
where $R_{t,M}=\sum_{i=1}^{M}R_{t+i}$, and $u,v$ are two complex numbers denoted by $u=u_{R}+is$ and $v=v_{R}+is$, with $u_{R}<0$, $v_{R}>0$ and $s\in\mathbb{R}$. The function, $g_{t,M}\left(u\right)=\mathbb{E}_{t}^{\mathbb{Q}}\left[\exp\left(uR_{t,M}\right)\right]$, is the conditional moment-generating function of cumulated returns over $M$ periods.
@@
Table 1 shows that the DHNG option pricing formula of Theorem 3 is far more accurate than those of Auxiliary and CHNG. This is true across all types of options and initial values of $\eta_{0}$. The absolute approximation error (MAE) is just 0.08% on average over all designs in Table 3, and the average error (ME) is -0.05%. Unsurprisingly, CHNG is the least accurate option pricing model. The assumptions that $\eta_{t}$ is constant leads to underpricing of options when $\eta_{0}$ is small and overpricing when $\eta_{0}$ is large. The option pricing formula for the Auxiliary structure systematically underprices options, and we observe that the pricing error is increasing in DTM. This is consistent with the results in Figure 4, where the predicted values of $h_{t+1}^{\ast},\ldots,h_{t+M}^{\ast}$ for the Auxiliary structure are downward-biased and increasingly so as $M$ increases. This has implications for empirical estimation of the Auxiliary structure. To rationalize the observed option prices, the Auxiliary structure will need to exaggerate the predicted values of $\log\eta_{t}$, such that Auxiliary is expected to yield larger predictions of $\log\eta_{t}$ than those based on DHNG when estimated with option prices. This is indeed what we find in our empirical analysis, see Table B.3 in the Appendix.
@@
Note: This table reports the approximation errors, $e_{i}=\left[\log{\rm IV}_{i}^{{\rm model}}-\log{\rm IV}_{i}^{{\rm true}}\right]\times 100$ for options with different Moneyness (Delta), days to maturity (DTM), and initial value of $\eta$. The true option prices are obtained with 10 million simulated paths from the true model, and the implied volatilities are computed using Black-Scholes formula. The column-wise averages for the mean absolute error (MAE) and the mean error (ME), are presented in the last two rows. CHNG, Auxiliary, and DHNG correspond to zero-, first-, and second-order approximations, respectively.
@@ 4 Observation-Driven Model for Variance Risk Aversion
In this section, we develop an observation-driven model for $\eta_{t}$, which facilitates an implementation of the results from the previous section. In practice, this requires inferring the current value of $\eta_{t}$ and estimate its dynamic model. To this end, we adopt an intuitive observation-driven model for $\eta_{t}$, inspired by the score-driven framework of Creal et al., (2013). We use the first-order conditions for minimizing pricing errors to define innovations to $\eta_{t}$, ensuring that $\eta_{t}$ is adjusted to reduce pricing errors on average. [^9] Alternatively, one could adopt state space approach. This is computationally more complicated without necessarily providing any benefit. Even when the true model is a state space model, the score-driven models are typically found to be competitive, see Koopman et al., (2016). The first-order condition of the log-likelihood function is presented in the empirical section. For now, it suffices to express the log-likelihood function in its generic form, $\sum_{t=1}^{T}\ell(R_{t},X_{t}|\mathcal{F}_{t-1})$, where $T$ is the sample size, and we should emphasize that it relies on information from both $\mathbb{P}$ and $\mathbb{Q}$. We factorize $\ell(R_{t},X_{t}|\mathcal{F}_{t-1})=\ell(R_{t}|\mathcal{F}_{t-1})+\ell(X_{t}|R_{t},\mathcal{F}_{t-1})$, where the latter can be expressed as $\ell_{t}(X_{t}|\mathcal{G}_{t})$. This likelihood term measures how well the observed derivative prices are explained by the statistical model for returns and the pricing kernel.
@@
$s_{t}=\frac{\nabla_{t}}{\sqrt{\mathbb{E}^{\mathbb{P}}\left[\nabla_{t}^{2}|\mathcal{G}_{t}\right]}},\quad\text{with}\quad\nabla_{t}=\frac{\partial\ell(X_{t}|\mathcal{G}_{t})}{\partial\log\eta_{t}}.$ (18)
When the score is evaluated at the true parameters we have
$\mathbb{E}^{\mathbb{P}}[s_{t}|\mathcal{G}_{t}]=0,\quad\text{and}\qquad\mathrm{var}^{\mathbb{P}}[s_{t}|\mathcal{G}_{t}]=1,$
such that $\mathbb{E}^{\mathbb{P}}[\varepsilon_{t}|\mathcal{G}_{t-1}]=0$ and $\mathrm{var}^{\mathbb{P}}[\varepsilon_{t}|\mathcal{G}_{t-1}]=\sigma^{2}$, since $s_{t}$ is $\mathcal{F}_{t}$-measurable. Note that this definition of the score does not ensure that $\{s_{t}\}$ is a sequence of iid random variables under $\mathbb{Q}$. This requires additional assumptions about the pricing errors. For instance, our assumption for pricing errors in Assumption 4 implies $s_{t}\sim iid\ N(0,1)$ under both in $\mathbb{P}$ and $\mathbb{Q}$ measures, see Theorem 5. From (12) and the specification (17) it follows that $\eta_{t}$ is $\mathcal{F}_{t-1}$-measurable, which ensures that information about derivatives prices are not used to price themselves.
@@ 4.1 The Log-Likelihood Function and the Form of Score
$\ell(R_{1},\ldots,R_{T},X_{1},\ldots,X_{T},\mathcal{F}_{0})=\sum_{t=1}^{T}\ell(R_{t},X_{t}|\mathcal{F}_{t-1})=\sum_{t=1}^{T}\ell(R_{t}|\mathcal{F}_{t-1})+\ell(X_{t}|R_{t},\mathcal{F}_{t-1}).$ (19)
This decomposition was also used by Christoffersen et al., (2013). The log-likelihood function for returns, $\ell(R_{t}|\mathcal{F}_{t-1})$, is that of the HNG model,
@@
which assumes that $z_{t}\sim iid\ N(0,1)$. To complete the model, we need to specify the log-likelihood function for the vector of derivative prices, $\ell(X_{t}|R_{t},\mathcal{F}_{t-1})=\ell(X_{t}|\mathcal{G}_{t})$. To this end, we let $X_{t}^{m}\in\mathcal{G}_{t}$ denote the vector of model-based prices and make the following assumption for pricing errors.
@@
Given the $\mathcal{G}_{t}$-filtration, the derivative pricing errors follow an iid multivariate normal distribution
$e_{t}=X_{t}-X_{t}^{m}\sim iid\ N\left(0,\sigma_{e}^{2}\Omega_{N_{t}}\right),$
with correlation matrix $\Omega_{N_{t}}$, where $N_{t}$ is the number of derivative prices at time $t$.
@@
The existing empirical studies that used derivative pricing errors in this manner, all assumed that the pricing errors were uncorrelated, i.e. $\Omega_{N_{t}}=I_{N_{t}}$. This is unrealistic, because contemporaneous pricing errors tend to be positively correlated. We will therefore allow for a non-zero common correlation, which is important for the statistical properties of the score, see Theorem 5 below. For instance, if we impose $\Omega_{N_{t}}=I_{N_{t}}$ in our empirical analysis with option prices, then it would result in a type of misspecification that induces an upward bias in the estimated variance of $s_{t}$.
@@
The correlation matrix for pricing errors has the form: $\Omega_{N_{t}}=\rho+(1-\rho)I_{N_{t}}$.
The correlation matrix in Assumption 5, $\Omega_{N_{t}}$, is known as an equicorrelation matrix. Having a common correlation for all pairs of pricing errors is particularly useful in applications where the panel of derivative prices is unbalanced over time. In this case, the dimension of $\Omega_{N_{t}}$ will be time-varying, but this is not problematic if an equicorrelation structure is assumed. A common correlation can be motivated by the assumption that pricing errors, $e_{i,t}=u_{t}+v_{i,t}$, have a common component, $u_{t}$, and uncorrelated idiosyncratic components, $v_{i,t}$, $i=1,\ldots,N_{t}$. This will bring about the structure in Assumption 5 with $\rho=\sigma_{u}^{2}/(\sigma_{u}^{2}+\sigma_{v}^{2})$ and $\sigma_{e}^{2}=\sigma_{u}^{2}+\sigma_{v}^{2}$.
@@
$\ell(X_{t}|\mathcal{G}_{t})=\ell(e_{t}|\mathcal{G}_{t})=-\tfrac{1}{2}\left(N_{t}\log(2\pi\sigma_{e}^{2})+\log\left|\Omega_{N_{t}}\right|+\sigma_{e}^{-2}e_{t}^{\prime}\Omega_{N_{t}}^{-1}e_{t}\right).$ (20)
The model for returns under $\mathbb{P}$ continues to be a time-homogeneous Heston-GARCH model, which does not depend on derivative prices. This is also true in the DHNG model, despite the time variation in the pricing kernel. For this reason, the model parameters can (if needed) be estimated by two-stage estimation, where the parameters in the GARCH model is estimated from returns in the first stage, followed by the remaining parameters being estimated in the second stage with derivative prices. [^11] In our empirical application we estimated the model by maximizing the full log-likelihood function, which was straightforward and did not cause computational issues. If needed, two-stage estimation can be adopted to reduce the computational burden, and the approach to estimation is not uncommon in this setting, see e.g. Broadie et al., (2007), Corsi et al., (2013), Christoffersen et al., (2013), and Majewski et al., (2015).
@@
which has zero conditional mean, $\mathbb{E}^{\mathbb{P}}[\nabla_{t}|\mathcal{G}_{t}]=0$, and conditional variance:
Moreover, the scaled score, $s_{t}\sim iid\ N(0,1)$, under both $\mathbb{P}$ and $\mathbb{Q}$ measures.
To obtain the score, we need to derive $\partial X_{t}^{m}/\partial\log\eta_{t}$, and this must be done separately for the VIX and option prices. These terms are derived in Appendix A.7.1 and Appendix A.7.2, respectively. Note that if $X_{t}^{m}$ is univariate (the case with a single derivative), then the scaled score simplifies to

### v4

@@ 2.2 Option Pricing: A Novel Analytical Approximation
The European call option price at time $t$ is given by the conditional risk-neutral expectation $C_{t}=e^{-r\left(T-t\right)}\mathbb{E}_{t}^{\mathbb{Q}}\left[\max\left(S_{T}-K,0\right)\right]$, where $T$ is the maturity date, $K$ is the strike price, and $S_{T}$ is the terminal price of underlying asset. Let $M=T-t$ be the number of periods to maturity, it is well-known that an affine structure of the MGF of future cumulative returns,
$g_{t,M}(s)=\mathbb{E}_{t}^{\mathbb{Q}}\left[\exp\left(s\sum_{i=1}^{M}R_{t+i}\right)\right],\quad s\in\mathbb{R}$
@@ 2.2.1 MGF with Predetermined Path
Suppose that Assumptions 1-2 hold and that $\log\eta_{t}$ take the predetermined path, $\bar{\boldsymbol{\eta}}_{t,M}$. Then the MGF for future cumulative returns under $\mathbb{Q}$ has the affine form,
@@ 2.2.2 Option Pricing with Stochastic Variance Risk Aversion
We now turn to the more challenging problem where $\boldsymbol{\eta}_{t,M}$ is stochastic. The lack of an affine structure thwarts the standard approach to obtaining closed-form option pricing formula, and it is common to resort to simulation methods in this case.
@@
Suppose that Assumptions 1-3 hold, then the approximate model-implied MGF, based on a second-order Taylor expansion, is given by
$\hat{g}_{t,M}(s)=g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M}^{e})\left[1+\frac{1}{2}{\rm tr}\{H_{t,M}(s)\Sigma_{M}\}\right],$
@@
Theorem 3 clarifies how random variation in $\eta_{t}$ impacts the MGF. [^8] The integrations in Theorem 3 can be numerically evaluated through the method in Appendix A.9. The core principle behind the proposed approximation method is akin to perturbation methods used to find approximate solutions to Schrödinger equations for which standard methods are not applicable. Perturbation methods also starts from an exact solution of a simpler, related problem, from which an approximate solution to the actual problem is deduced. Our approximation is based on a Taylor expansion about an $M$-dimensional vector, where the conditional moments of the terms in the expansion are deduced from Assumption 3. Next, we evaluate the accuracy of the approximation.
Figure 3: Risk-neutral densities of (six months) cumulative return, for different initial values of variance risk ratio, $\eta_{0}=0.70$ (left panel), $\eta_{0}=1.11$ (middle panel), and $\eta_{0}=1.70$ (right panel). The blue solid line represents the true density from 1 million simulations, the red dashed line is the (2nd-order) approximate density by DHNG, the gray dash-dotted line is the auxiliary model based on $\bar{\boldsymbol{\eta}}_{t,M}^{e}$ (1st-order), and the yellow dotted line is the (0th-order) density by CHNG with constant $\eta_{t}$. The model parameters are based on the estimates with option prices (the last column in Table 3).
Alt text: Risk-neutral densities for cumulative six-month returns, shown for three initial levels of the variance risk ratio: low, middle, and high. Each density is compared with three approximations—CHNG, the auxiliary model, and DHNG. DHNG provides the most accurate approximation.
@@ 2.3 Accuracy of Analytical Approximation
The option pricing formula in Theorem 3 involves a second-order approximation that accounts for the first two conditional moments of $\boldsymbol{\eta}_{t,M}$. [^9] If the distribution of $\boldsymbol{\varepsilon}_{t,M}$ is symmetric, (e.g. Gaussian), then it is an approximation to third-order. A first-order approximation is equivalent to assuming that $\eta_{t}$ will take the path of its conditional expectation, $\bar{\boldsymbol{\eta}}_{t,M}^{e}$. Thus, $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{e})$ is the MGF for the first-order approximation, which we previously labeled “Auxiliary”, because it was used as an intermediate step towards our preferred approximation, $\hat{g}_{t,M}(\cdot)$. Taking $\eta_{t}=\eta_{0}$ to be constant, $\bar{\boldsymbol{\eta}}_{t,M}^{c}=(\log\eta_{0},\ldots,\log\eta_{0})^{\prime}$, is a characteristic of CHNG, and this case can be interpreted as the zero-order “approximation”.
We first consider the risk-neutral density of cumulative returns. The blue solid lines in Figure 3 present the true risk-neutral densities for cumulative returns over six months for three initial values, $\eta_{0}=0.70$ (left panel), $\eta_{0}=1.11$ (middle panel), and $\eta_{0}=1.70$ (right panel), based on 1 million simulations. The data generation process is based on the structure in Theorem 1, using the parameter estimates from our empirical analysis of option prices with $\log\eta_{t}\sim\operatorname{AR}(1)$, normally distributed innovations, and $h_{1}$ initialized at its unconditional mean. In this design, the exponential of unconditional mean, $\exp\left(\mathbb{E}\log\eta_{t}\right)=\exp\left(\zeta\right)$, is used as the initial value $\eta_{0}$ in the middle panel.
The red dashed lines are the risk-neutral densities of DHNG implied by $\hat{g}_{t,M}(\cdot)$, the gray dot-dashed lines are based on the auxiliary, $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{e})$, and the yellow dotted line is the risk-neutral density for CHNG, i.e. $g_{t,M}(\cdot|\bar{\boldsymbol{\eta}}_{t,M}^{c})$. DHNG, which is based on the second-order approximation of Theorem 3, is accurate, whereas the Auxiliary structure fails to match the upper tail of the densities. The CHNG is the worst approximation of the true density in all cases. It is tied with Auxiliary in the middle panel, because the two are identical when $\log\eta_{0}$ is set to have its unconditional mean $\zeta$.
@@
Figure 4: Annualized variance, $\frac{252}{M}{\rm var}_{t}^{\mathbb{Q}}(\sum_{i=1}^{M}R_{t+i})\times 100\%$, skewness, and kurtosis of cumulative returns over $M$ days, are shown with three different initial variance risk ratios, $\eta_{0}=0.70$ (left panels), $\eta_{0}=1.11$ (middle panels), and $\eta_{0}=1.70$ (right panels). Blue lines represent true population moments, red dashed lines are those of DHNG, yellow dotted lines are those for CHNG, and the gray dash-dotted lines are those for the auxiliary “model”. Population moments were obtained from 1 million simulated trajectories, using a design based on the parameter estimates obtained with option prices (the last column in Table 3).
Alt text: Variance, skewness, and kurtosis of cumulative returns for three initial levels of the variance risk ratio: low, middle, and high. The figure compares the term structure of true moments with those estimated using CHNG, the auxiliary model, and DHNG, with DHNG providing the most accurate approximation.
@@
The conditional risk-neutral $k$-th moment for $k\in\mathbb{N}$ of the $M$-period ahead cumulative returns can be expressed as
@@
where $R_{t,M}=\sum_{i=1}^{M}R_{t+i}$, and $u,v$ are two complex numbers denoted by $u=u_{R}+is$ and $v=v_{R}+is$, with $u_{R}<0$, $v_{R}>0$ and $s\in\mathbb{R}$. The function, $g_{t,M}\left(u\right)=\mathbb{E}_{t}^{\mathbb{Q}}\left[\exp\left(uR_{t,M}\right)\right]$, is the conditional MGF of the cumulated returns over $M$ periods.
@@
Table 1 shows that the DHNG option pricing formula of Theorem 3 is far more accurate than those of Auxiliary and CHNG. This is true across all types of options and initial values of $\eta_{0}$. The absolute approximation error (MAE) is just 0.08% on average over all designs in Table 1, and the average error (ME) is -0.05%. Unsurprisingly, CHNG is the least accurate option pricing model. The assumptions that $\eta_{t}$ is constant leads to underpricing of options when $\eta_{0}$ is small and overpricing when $\eta_{0}$ is large. The option pricing formula for the Auxiliary structure systematically underprices options, and we observe that the pricing error is increasing in DTM. This is consistent with the results in Figure 4, where the predicted values of $h_{t+1}^{\ast},\ldots,h_{t+M}^{\ast}$ for the Auxiliary structure are downward-biased and increasingly so as $M$ increases. This has implications for empirical estimations under the Auxiliary structure. To rationalize the observed option prices, the Auxiliary structure will need to exaggerate the predicted values of $\log\eta_{t}$, such that Auxiliary is expected to yield larger predictions of $\log\eta_{t}$ than those based on DHNG when estimated with option prices. This is indeed what we find in the supplementary results, see Table S.3 in Appendix S.3.
@@
Note: This table reports the approximation errors, $e_{i}=\left[\log{\rm IV}_{i}^{{\rm model}}-\log{\rm IV}_{i}^{{\rm true}}\right]\times 100$ for options with different Moneyness (Delta), days to maturity (DTM), and initial value of $\eta$. The true option prices are obtained with 10 million simulated paths from the true model, and the implied volatilities are computed using Black-Scholes formula. The column-wise averages for the mean absolute error (MAE) and the mean error (ME), are presented in the last two rows. CHNG, Auxiliary, and DHNG correspond to zero-, first-, and second-order approximations, respectively. The model parameters are based on the empirical estimates obtained with option prices (the last column in Table 3).
@@ 3 Observation-Driven Model for Variance Risk Aversion
In this section, we develop an observation-driven model for $\eta_{t}$, which facilitates an implementation of the results from the previous section. In practice, this requires inferring the current value of $\eta_{t}$ and estimate its dynamic model. To this end, we adopt an intuitive observation-driven model for $\eta_{t}$, inspired by the score-driven framework of Creal et al., (2013). We use the first-order conditions for minimizing pricing errors to define innovations to $\eta_{t}$, ensuring that $\eta_{t}$ is adjusted to reduce pricing errors on average. [^10] Alternatively, one could adopt state space approach. This is computationally more complicated without necessarily providing any benefit. Even when the true model is a state space model, the score-driven models are typically found to be competitive, see Koopman et al., (2016). The first-order condition of the log-likelihood function is presented in the empirical section. For now, it suffices to express the log-likelihood function in its generic form, $\sum_{t=1}^{T}\ell(R_{t},X_{t}|\mathcal{F}_{t-1})$, where $T$ is the sample size, and we should emphasize that it relies on information from both $\mathbb{P}$ and $\mathbb{Q}$. We factorize $\ell(R_{t},X_{t}|\mathcal{F}_{t-1})=\ell(R_{t}|\mathcal{F}_{t-1})+\ell(X_{t}|R_{t},\mathcal{F}_{t-1})$, where the latter can be expressed as $\ell(X_{t}|\mathcal{G}_{t})$. This likelihood term measures how well the observed derivative prices are explained by the statistical model for returns and the pricing kernel.
@@
$s_{t}=\frac{\nabla_{t}}{\sqrt{\mathbb{E}_{t}^{\mathbb{P}}\left(\nabla_{t}^{2}\right)}},\quad\text{with}\quad\nabla_{t}=\frac{\partial\ell(X_{t}|\mathcal{G}_{t})}{\partial\log\eta_{t}}.$ (9)
Since $s_{t}$ is $\mathcal{F}_{t}$-measurable, we have $\varepsilon_{t+1}\in\mathcal{F}_{t}\subset\mathcal{G}_{t+1}$ as required by Assumption 3. Additionally, when the score is evaluated at the true parameters we have
$\mathbb{E}_{t}^{\mathbb{P}}\left(s_{t}\right)=0,\quad\text{and}\qquad\mathrm{var}_{t}^{\mathbb{P}}\left(s_{t}\right)=1,$
such that $\mathbb{E}_{t}^{\mathbb{P}}(\varepsilon_{t+1})=0$ and $\mathrm{var}_{t}^{\mathbb{P}}(\varepsilon_{t+1})=\sigma^{2}$. Note that this definition of the score does not ensure that $\{s_{t}\}$ is a sequence of iid random variables with zero mean and unit variance under $\mathbb{Q}$, as needed by Assumption 3. This requires additional distributional assumptions about the pricing errors. For instance, we will make assumptions about the pricing errors that imply $s_{t}|\mathcal{G}_{t}\sim iid\ N(0,1)$ under both in $\mathbb{P}$ and $\mathbb{Q}$ measures (see Theorem 5 below), in which case the score satisfies all the requirements in Assumption 3. From (6) and the specifications (8)-(9), it follows that our constructed $\eta_{t}$ is $\mathcal{F}_{t-1}$-measurable, which ensures that the information about current derivative prices are not used to price themselves.
@@ 3.1 The Log-Likelihood Function and the Form of Score
$\ell(R_{1},\ldots,R_{T},X_{1},\ldots,X_{T},\mathcal{F}_{0})=\sum_{t=1}^{T}\ell(R_{t},X_{t}|\mathcal{F}_{t-1})=\sum_{t=1}^{T}\ell(R_{t}|\mathcal{F}_{t-1})+\ell(X_{t}|R_{t},\mathcal{F}_{t-1}).$
This decomposition was also used by Christoffersen et al., (2013). The log-likelihood function for returns, $\ell(R_{t}|\mathcal{F}_{t-1})$, is that of the HNG model in (1)-(2),
@@
which assumes that $z_{t}\overset{\mathbb{P}}{\sim}iid\ N(0,1)$. To complete the model, we need to specify the log-likelihood function for the vector of derivative prices, $\ell(X_{t}|R_{t},\mathcal{F}_{t-1})=\ell(X_{t}|\mathcal{G}_{t})$. To this end, we let $X_{t}^{m}\in\mathcal{G}_{t}$ denote the vector of model-based derivative prices and make the following assumption for pricing errors.
@@
The derivative pricing errors follow an iid multivariate normal distribution
$e_{t}=X_{t}-X_{t}^{m}\left|\mathcal{G}_{t}\right.\overset{\mathbb{P}}{\sim}iid\ N\left(0,\sigma_{e}^{2}\Omega_{N_{t}}\right),$
with correlation matrix $\Omega_{N_{t}}$, where $N_{t}$ is the number of derivative prices at time $t$. Additionally, $e_{t}$ is independent of the return shock $z_{\tau}$ for all $\tau$.
@@
The existing empirical studies that used derivative pricing errors in this manner, all assumed that the pricing errors were uncorrelated, i.e. $\Omega_{N_{t}}=I_{N_{t}}$ where $I_{N_{t}}$ is the $N_{t}$-dimensional identity matrix. This is unrealistic, because contemporaneous pricing errors tend to be positively correlated. We will therefore allow for a non-zero common correlation, which is important for the statistical properties of the score, see Theorem 5 below. For instance, if we impose $\Omega_{N_{t}}=I_{N_{t}}$ in our empirical analysis with option prices, then it would result in a type of misspecification that induces an upward bias in the estimated variance of $s_{t}$.
@@
The correlation matrix for pricing errors has the form: $\Omega_{N_{t}}=(1-\rho)I_{N_{t}}+\rho U_{N_{t}}$, where $I_{N_{t}}\in\mathbb{R}^{N_{t}\times N_{t}}$ is the identity matrix and $U_{N_{t}}\in\mathbb{R}^{N_{t}\times N_{t}}$ is a matrix of ones.
The correlation matrix in Assumption 5, $\Omega_{N_{t}}$, is known as an equicorrelation matrix. Having a common correlation for all pairs of pricing errors is particularly useful in applications where the panel of derivative prices is unbalanced over time. In this case, the dimension of $\Omega_{N_{t}}$ will be time-varying, but this is not problematic if an equicorrelation structure is assumed. A common correlation can be motivated by the assumption that pricing errors, $e_{i,t}=u_{t}+v_{i,t}$, have a common component, $u_{t}$, and uncorrelated idiosyncratic components, $v_{i,t}$, $i=1,\ldots,N_{t}$. This will bring about the structure in Assumption 5 with $\rho=\sigma_{u}^{2}/\sigma_{e}^{2}$ and $\sigma_{e}^{2}=\sigma_{u}^{2}+\sigma_{v}^{2}$.
@@
$\ell(X_{t}|\mathcal{G}_{t})=\ell(e_{t}|\mathcal{G}_{t})=-\tfrac{1}{2}\left(N_{t}\log(2\pi\sigma_{e}^{2})+\log\left|\Omega_{N_{t}}\right|+\sigma_{e}^{-2}e_{t}^{\prime}\Omega_{N_{t}}^{-1}e_{t}\right).$
The model for returns under $\mathbb{P}$ continues to be a time-homogeneous Heston-GARCH model, which does not depend on derivative prices. This is also true in the DHNG model, despite the time variations in the pricing kernel. For this reason, the model parameters can (if needed) be estimated by a two-stage estimation method, where the parameters in the GARCH model is estimated from returns in the first stage, followed by the remaining parameters being estimated in the second stage with derivative prices. [^12] In our empirical application we estimated the model by maximizing the full log-likelihood function, which was straightforward and did not cause computational issues. If needed, two-stage estimation can be adopted to reduce the computational burden, and the approach to estimation is not uncommon in this setting, see e.g. Broadie et al., (2007), Corsi et al., (2013), Christoffersen et al., (2013), and Majewski et al., (2015).
@@
which has zero conditional mean, $\mathbb{E}_{t}^{\mathbb{P}}\left(\nabla_{t}\right)=0$, and conditional variance:
$\mathbb{E}_{t}^{\mathbb{P}}\left(\nabla_{t}^{2}\right)=\frac{1}{\sigma_{e}^{2}}\left(\frac{\partial X_{t}^{m}}{\partial\log\eta_{t}}\right)^{\prime}\Omega_{N_{t}}^{-1}\left(\frac{\partial X_{t}^{m}}{\partial\log\eta_{t}}\right).$
Moreover, the scaled score is $\mathcal{F}_{t}$-measurable and satisfies $s_{t}|\mathcal{G}_{t}\sim iid\ N(0,1)$, under $\mathbb{P}$ and $\mathbb{Q}$ measures.
To obtain the score, we need to derive $\partial X_{t}^{m}/\partial\log\eta_{t}$, and this must be done separately for the VIX and option prices. These terms are derived in Appendix A.8. Note that if $X_{t}^{m}$ is univariate (the case with a single derivative), then the scaled score simplifies to